# Unveiling AI’s Blind Spots: An Oracle for In-Domain, Out-of-Domain, and Adversarial Errors

### v2

@@ 4.2 Mentor architectures matter in error predictions
To computationally model the decision boundary of a mentee using a mentor, the mentor requires more complex architectures with a larger number of parameters than the mentee. Indeed, from Fig. 3, over all the datasets, we observed that utilizing ViT (V) as the mentor backbone consistently achieves higher accuracy across all error types of ViT-based and ResNet-based mentees compared to the mentor based on ResNet50 (R). One example of this performance disparity is observed in the context of the adversarial attack error type for CIFAR-10. The ViT-based mentor attains an accuracy of 74.95%, substantially higher than the accuracy of 63.99% for the ResNet-based mentor.
Loss landscape analysis. The performance difference between mentors’ architectures is due to ViT’s superior ability to identify features from error patterns. Its self-attention mechanism captures complex relationships among data samples, providing a deeper understanding of the mentee’s loss landscape, particularly in modelling irregular, rugged landscapes with sharp peaks and valleys.
ID | SpN | GaB | Spat | Sat | PGD | CW | Jitter | PIFGSM | Average | |
SER | 50.2 | 49.8 | 49.9 | 50.1 | 49.9 | 50.1 | 49.9 | 49.9 | 49.9 | 50.0 |
MCP () | 77.3 | 76.8 | 77.4 | 78.3 | 76.6 | 71.8 | 72.9 | 71.7 | 69.4 | 74.7 |
CPE () | 78.0 | 75.2 | 75.9 | 78.0 | 76.2 | 72.2 | 72.8 | 72.7 | 69.3 | 74.5 |
DTC () | 52.4 | 51.2 | 52.9 | 52.7 | 52.5 | 51.4 | 51.6 | 51.7 | 51.6 | 52.0 |
ConfidNet (Corbière et al., 2019) | 67.3 | 65.9 | 68.5 | 70.8 | 67.9 | 68.8 | 69.0 | 70.1 | 67.1 | 68.4 |
TrustScore (Jiang et al., 2018) | 70.0 | 67.5 | 71.1 | 73.4 | 71.3 | 73.2 | 72.5 | 73.9 | 72.7 | 71.7 |
SSL (Luo et al., 2021) | 66.3 | 66.7 | 69.7 | 71.9 | 70.0 | 69.1 | 70.1 | 70.2 | 66.7 | 69.0 |
SuperMentor (ours) | 78.9 | 73.6 | 74.6 | 78.3 | 73.6 | 83.0 | 78.4 | 79.9 | 72.3 | 77.0 |
@@ 4.3 Training on images with smaller perturbations helps error predictions
Although adversarial images have been demonstrated to aid in error prediction (Sec. 4.1), it remains unclear whether adversarial images with varying degrees of image distortion exhibit the same effect. A straightforward method to regulate the level of image distortion caused by adversarial attacks is to set the perturbation bound . We employ four corruption levels by setting , , , and . We use the adversarial attack PIFGSM as an example since the error patterns from PIFGSM are most effective for the mentor’s prediction (see Appendix, Fig. S3- S5). As shown in Fig. 5, the mentor’s accuracy significantly decreases as the distortion level increases. In particular, for the C10-AA-PIFGSM, the accuracy at level 1 is , which is notably higher than at level 4. Our findings suggest that adversarial attacks employing smaller perturbations yield more benefits for mentor error prediction. This phenomenon can be attributed to the fact that adversarial images with minimal perturbations maintain closer proximity to the decision boundary of a mentee.
@@ 4.4 Mentors generalize across mentees
In Sec. 4.1, mentors have demonstrated their ability to learn the error patterns of mentees. This observation raises an important question: can the error patterns learned from one mentee (mentee A) be generalized to another mentee (mentee B) when the two mentees employ different model architectures? To explore this, we evaluate all 324 mentors, whose performances are depicted in Appendix, Tab. S6- S8, on the alternate mentee. Specifically, mentors trained on the errors of the ResNet50 mentee are tested on the predictions of the ViT mentee, and vice versa. The outcomes of these evaluations are illustrated in Fig. 5. Surprisingly, most points lie near the dashed diagonal line, implying that the mentors’ performance does not significantly deteriorate when evaluated on the predictions of different mentee architectures. This finding indicates that ResNet50 and ViT mentees tend to produce similar error patterns when trained on the same dataset.
4.5 Our SuperMentor outperforms all baselines
@@
We demonstrate the effectiveness of SuperMentor by comparing it with baselines introduced in Sec. 3.4. The results are shown in Tab. 1. It can be seen that our SuperMentor consistently achieves higher average accuracies over all error types compared to all baselines. Notably, in the AA scenarios, SuperMentor has significantly higher accuracy than the baselines. For example, SuperMentor achieves an error prediction accuracy of 83.0% on IN-AA-PGD of the ResNet50 mentee, whereas the best baseline only reaches 73.3%.
@@
Baseline methods like SER, MCP, CPE, and DTC rely on fixed thresholds or predetermined values of manually defined criteria such as confidence or entropy, making them less adaptable to OOD and AA scenarios. In contrast, SuperMentor leverages deep neural networks to capture the complexity of error prediction. Additionally, methods like ConfidNet, TrustScore, and SSL are trained on ID error types. As shown in Tab. 1, these baselines are inferior to SuperMentor, which is trained exclusively on AA data. This suggests that the source of error data used to train the mentor plays a more critical role than the choice of loss functions or the sophisticated training strategies in these baselines.
We conduct extra network analysis of our SuperMentor. Specifically, we include an ablation study on the distillation loss (Appendix, Sec. F). The removal of leads to a significant drop in Accuracy, highlighting the importance of . Moreover, we provide the visualization of SuperMentor’s embeddings based on the correctness of a mentee’s prediction across three error types (Appendix, Sec. G). Two distinct clusters are observed between samples wrongly or correctly classified by the mentee. This suggests that our SuperMentor is capable of accurately predicting the errors of a mentee regardless of the error sources.
@@ 5 Discussion and Conclusion
Despite the promising results, our framework may raise concerns regarding the vulnerabilities of AI mentors. Indeed, vulnerability is a well-known challenge for all Deep Neural Network (DNN)-based systems, yet this has not hindered their transformative applications in the real world. For example, DNNs are used to detect diabetic retinopathy (Jumper et al., 2021; Alyoubi et al., 2020), outperforming traditional diagnostic methods. Similarly, our work uses one AI to diagnose the other AI, which is a valid and impactful use case. To demonstrate its real-world utility, we extended our experiments to a medical image classification task for colorectal cancer diagnosis (Appendix, Sec. H). Our mentor achieves high error prediction accuracy compared to all competitive baselines, showcasing its practical uses in high-stake applications. More importantly, beyond the OOD domains evaluated in Sec. 4.5, results in Appendix, Sec. I show that our SuperMentor remains robust to mentee error patterns across an additional set of five unseen OOD domains. Moreover, the vulnerabilities of the mentor AI do not overlap with those of the mentees since mentors operate independently of the mentee’s architecture or training details, reducing the risk of shared weaknesses. To validate this point, we use EigenCAM to visualize the behavior of both our ViT SuperMentor and ViT mentee on example images (Appendix, Fig. S2). From Fig. S2 (a), the mentor does not merely replicate the mentee’s learning patterns, as their activation maps for identical input images can differ significantly. Interestingly, our AI mentor can even serve as a valuable diagnostic tool for the AI mentees. In Appendix, Fig. S2 (b), the mentor highlights the key feature pattern diagnostic for object recognition that the mentee tends to overlook. This suggests that the mentor can identify vulnerabilities in DNN-based models.
@@
Appendix A Dataset split for each error source used in mentor training
As mentioned in Sec. 3.3, for any given dataset of a mentor, let and represent the sets of correctly and incorrectly classified images by a mentee. The sizes of and can vary significantly, depending on the mentee’s classification performance. A mentee with high recognition accuracy will have more correct classifications (big ) and fewer incorrect ones (small ). To create a balanced test set for a mentor, we select equal numbers of correctly and incorrectly classified samples. The remaining samples are used for training. The details of the dataset split are shown in Appendix, Tab. S1. To address the long-tail problem in the training set, during each training epoch for a mentor, we randomly generate a batch of samples that includes an equal number of correctly and incorrectly classified images by the mentee.
Mentee | Error Source | CIFAR-10 | CIFAR-100 | ImageNet-1K | ||||
ResNet50 | ID | 151/9547 | 151/151 | 773/7681 | 773/773 | 5967/32099 | 5967/5967 | |
OOD | SpN | 690/7930 | 690/690 | 1889/4333 | 1889/1889 | 9984/20048 | 9984/9984 | |
GaB | 149/9553 | 149/149 | 760/7720 | 760/760 | 8013/25963 | 8012/8012 | ||
Spat | 222/9336 | 221/221 | 990/7032 | 989/989 | 7042/28874 | 7042/7042 | ||
Sat | 240/9282 | 239/239 | 1309/6072 | 1309/1309 | 8187/25439 | 8187/8187 | ||
AA | Jitter | 338/8988 | 337/337 | 1054/6840 | 1053/1053 | 7591/27227 | 7591/7591 | |
PGD | 447/8661 | 446/446 | 1180/6460 | 1180/1180 | 9009/22973 | 9009/9009 | ||
CW | 487/8539 | 487/487 | 1120/6642 | 1119/1119 | 8102/25694 | 8102/8102 | ||
PIFGSM | 1613/5161 | 1613/1613 | 2090/3732 | 2089/2089 | 11226/16322 | 11226/11226 | ||
ViT | ID | 128/9618 | 127/127 | 675/7977 | 674/674 | 4733/35801 | 4733/4733 | |
OOD | SpN | 286/9144 | 285/285 | 1155/6535 | 1155/1155 | 6019/31945 | 6018/6018 | |
GaB | 130/9610 | 130/130 | 678/7966 | 678/678 | 6402/30794 | 6402/6402 | ||
Spat | 170/9490 | 170/170 | 809/7573 | 809/809 | 5351/33947 | 5351/5351 | ||
Sat | 227/9319 | 227/227 | 1219/6345 | 1218/1218 | 5883/32351 | 5883/5883 | ||
AA | Jitter | 552/8344 | 552/552 | 1232/6304 | 1232/1232 | 10325/19025 | 10325/10325 | |
PGD | 649/8053 | 649/649 | 1410/5770 | 1410/1410 | 14960/11680 | 11680/11680 | ||
CW | 446/8664 | 445/445 | 1136/6592 | 1136/1136 | 8614/24158 | 8614/8614 | ||
PIFGSM | 799/7605 | 798/798 | 1812/4564 | 1812/1812 | 15038/11654 | 11654/11654 |
Appendix B Details of baselines
@@
Maximum Class Probability (MCP) (Hendrycks & Gimpel, 2016): As mentioned in Sec. 3.1, the mentee’s output logit is denoted as . The maximum class probability of the mentee’s output is defined as where is the softmax function. The MCP indicates how confident the mentee is in its most probable prediction. We set a threshold to distinguish between the mentee’s confident and unconfident predictions. Specifically, if , the mentee’s prediction will be considered a correct prediction due to its high confidence; otherwise, it will be regarded as an incorrect prediction. We employ three predefined thresholds for : 0.5, 0.7, and 0.9.
@@
Appendix C Example of calculating a mentor’s average accuracy using the proposed evaluation metric
As an example of computing the accuracy of a mentor evaluated using the metric presented in Sec. 3.4, we consider a mentor trained on C10-AA-PIFGSM and evaluate it on all the following datasets with their accuracies of 10% on C10-ID, 20% on C10-OOD-SpN, 30% on C10-OOD-GaB, 40% on C10-OOD-Spat, 50% on C10-OOD-Sat, 60% on C10-AA-Jitter, 70% on C10-AA-PGD, 80% on C10-AA-CW, and 90% on C10-AA-PIFGSM testing samples (See Appendix, Tab. S1 for testing splits), the average accuracy of this mentor is calculated as (10% + 20% + 30% + 40% + 50% + 60% + 70% + 80% + 90%) / 9 = 50%. For simplicity, we refer to this average accuracy across all nine error sources as Accuracy. A mentor randomly guessing whether a mentee’s image classification is correct or incorrect for a given image would achieve an accuracy of 50%.
Appendix D Joint training
As mentioned in Sec. 4.1, we add an experiment by training a Vision Transformer (ViT)-based mentor model on the correctness of a mixture of ID, OOD, and AA data for the ResNet-based mentee model from the CIFAR-10 dataset. The OOD data was corrupted with speckle noise, and the AA data was generated using PIFGSM. Compared to training the mentor solely on AA data which achieved an average accuracy of 78.0%, including ID and OOD data improved the average accuracy slightly to 79.1%. This minor improvement suggests a significant overlap of mentee error patterns among ID, OOD, and AA data. In other words, training mentors on OOD or ID data do not necessarily benefit mentors. Therefore, considering the associated computational costs, it is recommended to train mentors exclusively on AA data in practical applications. This finding underscores the significance of our study by providing critical insights into the optimal training practices for mentor models.

### v3

@@ 4.2 Mentor architectures matter in error predictions
To computationally model the decision boundary of a mentee using a mentor, the mentor requires more complex architectures with a larger number of parameters than the mentee. Indeed, from Fig. 3, over all the datasets, we observed that utilizing ViT (V) as the mentor backbone consistently achieves higher accuracy across all error types of ViT-based and ResNet-based mentees compared to the mentor based on ResNet50 (R). One example of this performance disparity is observed in the context of the adversarial attack error type on C10. The mentors leveraging a ViT backbone achieves 73.1% accuracy in predicting the correctness of ResNet50 mentee’s outputs, which is significantly higher than the 61.9% accuracy of the ResNet-based mentor (compare R-R versus R-V).
Loss landscape analysis. The performance difference between mentors’ architectures is due to ViT’s superior ability to identify features from error patterns. Its self-attention mechanism captures complex relationships among data samples, providing a deeper understanding of the mentee’s loss landscape, particularly in modelling irregular, rugged landscapes with sharp peaks and valleys. To validate this point, we include loss landscape analyses in Appendix, Sec. D. Results show that ViT mentors possess a notably broader loss landscape than ResNet50 mentors.
@@ 4.3 Training on images with smaller perturbations helps error predictions
Although adversarial images have been demonstrated to aid in error prediction (Sec. 4.1), it remains unclear whether adversarial images with varying degrees of image distortion exhibit the same effect. A straightforward method to regulate the level of image distortion caused by adversarial attacks is to set the perturbation bound . We employ four corruption levels by setting , , , and . We use the adversarial attack PIFGSM as an example since the error patterns from PIFGSM are most effective for the mentor’s prediction (see Appendix, Fig. S4- S6). As shown in Fig. 5, the mentor’s accuracy significantly decreases as the distortion level increases. In particular, for the C10-AA-PIFGSM, the accuracy at level 1 is , which is notably higher than at level 4. Our findings suggest that adversarial attacks employing smaller perturbations yield more benefits for mentor error prediction. This phenomenon can be attributed to the fact that adversarial images with minimal perturbations maintain closer proximity to the decision boundary of a mentee.
@@ 4.4 Mentors generalize across mentees
In Sec. 4.1, mentors have demonstrated their ability to learn the error patterns of mentees. This observation raises an important question: can the error patterns learned from one mentee (mentee A) be generalized to another mentee (mentee B) when the two mentees employ different model architectures? To explore this, we evaluate all 324 mentors on the alternate mentee. Specifically, mentors trained on the errors of the ResNet50 mentee are tested on the predictions of the ViT mentee, and vice versa. The outcomes of these evaluations are illustrated in Fig. 5. Surprisingly, most points lie near the dashed diagonal line, implying that the mentors’ performance does not significantly deteriorate when evaluated on the predictions of different mentee architectures. This finding indicates that ResNet50 and ViT mentees tend to produce similar error patterns when trained on the same dataset.
In addition, we evaluate the mentors’ generalization ability on natural adversarial samples introduced in (Hendrycks et al., 2021b), where each image is misclassified by multiple mentees simultaneously. See details in Appendix, Sec. F. Results show that, beyond natural adversarial examples, the mentors also exhibit robust generalization on non-natural adversarial examples.
4.5 Analysis on our SuperMentor reveals key insights
@@
We demonstrate the effectiveness of SuperMentor by comparing it with the baselines introduced in Sec. 3.4. To conduct comprehensive evaluations, we extend image perturbations from different error sources to multiple severity levels. Specifically, we use corruption levels 1, 3, and 5 for SpN, GaB, Spat, and Sat error sources as specified in (Hendrycks & Dietterich, 2019); set to , , and for PIFGSM, Jitter, and PGD error sources; and vary the learning rate as 0.01, 0.1, and 1.0 for the CW error source.
@@
As shown in Fig. 6(a), our SuperMentor can achieve higher average accuracies across all severity levels and error types than baselines. Notably, as indicated in Appendix, Tab. S1, in the AA scenarios, SuperMentor has significantly higher accuracy than the baselines. For example, SuperMentor achieves an error prediction accuracy of 73.1% on IN-AA-PGD of the ResNet50 mentee, whereas the best baseline only reaches 64.7%. Baseline methods like SER, MCP, CPE, and DTC rely on fixed thresholds or predetermined values of manually defined criteria such as confidence or entropy, making them less adaptable to AA scenarios. In contrast, SuperMentor leverages deep neural networks to capture the complexity of error prediction. Additionally, methods like ConfidNet, TrustScore, and SSL are trained on ID error types. As shown in Fig. 6(a), these baselines are inferior to SuperMentor, which is trained exclusively on AA data. This suggests that the source of error data used to train the mentor plays a more critical role than the choice of loss functions or the sophisticated training strategies in these baselines. Moreover, we provide the visualization of SuperMentor’s embeddings based on the correctness of a mentee’s prediction across three error types (Appendix, Sec. G). Two distinct clusters are observed between samples wrongly or correctly classified by the mentee. This suggests that our SuperMentor is capable of accurately predicting the errors of a mentee regardless of the error sources.
Next, we examine the effect of the distillation loss (Fig. 2) on the mentors’ performance. The results are presented in Fig. 6(b). It is clear that excluding results in a decrease in mentors’ accuracy across all datasets. For example, in the C10 dataset, the average accuracy of mentors decreases from to . This suggests that encourages mentors to learn the fine-grained decision boundaries among different object classes of a mentee. Alternatively, instead of utilizing the mentee’s logits, mentors can incorporate a cross-entropy loss to align the mentor’s predicted object class labels with those of the mentee, denoted as . From Fig. 6(b), we observe that replacing with leads to a drop in accuracy. This is due to the fact that the mentee’s logits contain more information than its class labels.
Finally, we provide EigenCAM (Gildenblat & contributors, 2021; Muhammad & Yeasin, 2020) visualization in Fig. 6(c) to illustrate the following two points: 1) The vulnerabilities of the mentee do not overlap with those of the mentors. As illustrated in Fig. 6(c), the mentor does not merely replicate the mentee’s learning patterns, as their activation maps for identical input images can differ significantly. For example, in the brambling images shown in Fig. 6(c), the activation maps of the mentor and mentee diverge greatly when the input image is ID or subjected to AA using PIFGSM method. In the IN-AA-PIFGSM image, the mentee demonstrates vulnerability under adversarial attacks, whereas our SuperMentor does not exhibit such vulnerabilities on this AA image, consistently focusing on the brambling object. This indicates the vulnerabilities of mentors and mentees are distinct due to their non-overlapping objectives. 2) Our AI mentor can serve as a valuable diagnostic tool for AI mentees. As supported by the baseball images in Fig. 6(c), although the mentee correctly classifies the ID image, its activation map does not concentrate on the baseball but instead highlights background cues (spurious features). When the input image is subjected to IN-AA-PIFGSM attacks, the mentee fails to make accurate predictions. In contrast, our SuperMentor effectively focuses on the stitches of the baseball regardless of whether the image is corrupted or adversarially attacked. This suggests that the mentor can identify vulnerabilities in the mentee, even though they both share the ViT architecture.
@@ 5 Discussion and Conclusion
Despite the promising results, our framework may raise concerns regarding the vulnerabilities of AI mentors. Indeed, vulnerability is a well-known challenge for all deep neural networks (DNNs), yet this has not hindered their transformative applications in the real world. For example, DNNs are used to detect diabetic retinopathy (Jumper et al., 2021; Alyoubi et al., 2020), outperforming traditional diagnostic methods. Similarly, our work uses one AI to diagnose the other AI, which is a valid and impactful use case. To demonstrate its real-world utility, we extended our experiments to a medical image classification task for colorectal cancer diagnosis (Appendix, Sec. H). Our mentor achieves high error prediction accuracy compared to all competitive baselines, showcasing its practical uses in high-stake applications. More importantly, beyond the OOD domains evaluated in Sec. 4.5, results in Appendix, Sec. I show that our SuperMentor remains robust to mentee error patterns across an extra set of five unseen OOD domains. Moreover, the vulnerabilities of the mentor AI do not overlap with those of the mentees since mentors operate independently of the mentee’s architecture or training details, reducing the risk of shared weaknesses (see Sec.4.5).
@@
Acknowledgements
This research is supported by the National Research Foundation, Singapore under its NRFF award NRF-NRFF15-2023-0001 and Mengmi Zhang’s Startup Grant from Nanyang Technological University.
@@ Acknowledgements
Appendix A Details of baselines
@@
Maximum Class Probability (MCP) (Hendrycks & Gimpel, 2017): As mentioned in Sec. 3.1, the mentee’s output logit is denoted as . The maximum class probability of the mentee’s output is defined as where is the softmax function. The MCP indicates how confident the mentee is in its most probable prediction. We set a threshold to distinguish between the mentee’s confident and unconfident predictions. Specifically, if , the mentee’s prediction will be considered a correct prediction due to its high confidence; otherwise, it will be regarded as an incorrect prediction. We employ three predefined thresholds for : 0.5, 0.7, and 0.9.
@@
Appendix B Example of calculating a mentor’s average accuracy using the proposed evaluation metric
As an example of computing the accuracy of a mentor evaluated using the metric presented in Sec. 3.4, we consider a mentor trained on C10-AA-PIFGSM and evaluate it on all the following datasets with their accuracies of 10% on C10-ID, 20% on C10-OOD-SpN, 30% on C10-OOD-GaB, 40% on C10-OOD-Spat, 50% on C10-OOD-Sat, 60% on C10-AA-Jitter, 70% on C10-AA-PGD, 80% on C10-AA-CW, and 90% on C10-AA-PIFGSM testing samples, the average accuracy of this mentor is calculated as (10% + 20% + 30% + 40% + 50% + 60% + 70% + 80% + 90%) / 9 = 50%. For simplicity, we refer to this average accuracy across all nine error sources as Accuracy. A mentor randomly guessing whether a mentee’s image classification is correct or incorrect for a given image would achieve an accuracy of 50%.
Appendix C Joint training
As mentioned in Sec. 4.1, we add an experiment by training a Vision Transformer (ViT)-based mentor model on the correctness of a mixture of ID, OOD, and AA data for the ResNet-based mentee model from the C10 dataset. The OOD data was corrupted with speckle noise (SpN), and the AA data was generated using PIFGSM. Compared to training the mentor solely on PIFGSM data which achieved an average accuracy of 74.3%, including ID and SpN data improved the average accuracy slightly to 77.5%. However, this joint training setup involves 21,000 samples—three times more than the 7,000 used for PIFGSM-only training. The marginal improvement observed in joint training may be attributed to the larger training sample size. This suggests that the quantity of samples is not the key factor in enhancing a mentor’s error prediction performance. Training mentors on samples that accurately capture the error patterns of mentees is crucial.
In addition, we introduce another mentor that is jointly trained on all error sources (ID, OOD, and AA, a total of 9 error sources) from the C10 dataset, achieving an average evaluation accuracy of 78.4%. In this scenario, all testing samples for the mentor are in distribution because the mentor has seen samples from every error source during training. Despite this, the improvement compared to a mentor trained solely on the PIFGSM error source (78.4% vs. 74.3%) is marginal. Consequently, this finding further indicates: (1) Training mentors exclusively on the AA error source offers a more data-efficient strategy for real-world applications. (2) Remarkably, a mentor trained solely on AA generalizes well to unseen error sources, achieving performance comparable to the upper bound obtained by jointly training on all error sources.
Appendix D Loss landscape analysis
We conducted two quantitative loss landscape analyses using the method from (Li et al., 2018) to support the claims in Sec. 4.1 and Sec. 4.2. In these analyses, we apply small perturbations to the mentor’s weights and observe the resulting changes in the mentors’ average accuracy.
As shown in Appendix, Fig. S2, we examine mentors trained on three different error types: ID, OOD, and AA from the C100 dataset. The results indicate that mentors trained on AA errors exhibit a wider loss landscape than those trained on ID or OOD errors. This suggests that mentors trained with adversarial images capture the generic features for predicting the mentee’s decision-making. This finding strongly supports our loss landscape analysis discussed in Sec. 4.1.
The second loss landscape analysis in Appendix, Fig. S2 examines the mentor’s error prediction performance using two different backbones: ResNet50 and ViT. The figure shows that mentors with ViT architectures have a considerably wider loss landscape compared to those with ResNet50. This further reinforces our claims in Sec. 4.2 that transformer-based mentor models outperform their ResNet-based counterparts in error prediction.